# A Longitudinal Study of Child Wellbeing Assessment via Online Interactions with a Social Robot

### v2

@@ 5 Results
To sum up, our results show that across all three sessions, RCADS responses have been significantly higher in the high tertile than in the low tertile. RCADS responses have also been significantly higher for the high tertile when comparing between modes of questionnaire administration (robot-report, self-report and parent-report). Longitudinally, robot anxiety measurement has been seen to decrease for the overall population. Further, the physical presence (remote vs. in-lab) of the robot does affect users’ anxiety attributed to the robot (robot anxiety increases for the low and med tertile but decreases for the high tertile) and the perceived intelligence of the robot (decreases across all tertiles). Across genders, we have observed that the RCADS responses were significantly higher for the high tertile for both boys and girls in session 1 and only for boys in session 2. We also observed that RCADS responses to the robot were significantly more negative for boys as compared to their parent’s perception of their wellbeing. Finally, girls have a higher perception of the robot as their confidante as compared with boys and the physical presence of the robot does affect this perception.
6 Discussion
@@
As seen from the construct validity analyses mentioned in Section 4.2, we employed a confirmatory PCA to explore the underlying structure of the SMFQ and the RCADS. We found that both scales predominantly revolve around a single principal component, suggesting a consistent underlying factor in each questionnaire explaining high variance of childrens’ well being. It is important to note that not all items in these questionnaires contributed equally to this primary factor. Specifically, some items in both the SMFQ and RCADS did not significantly load onto the principal component, indicating that they might not align as closely with the core factor that these scales measure within CRI. It should be acknowledged that RCADS (as used in this study) is aimed at assessing the concepts of generalised anxiety, panic disorder, and low mood [see 84], thereby potentially affecting factor loadings. However, the single factor identified evidence of comorbidities between low mood and anxiety symptoms, which are normally highly related [84].
@@
Nevertheless, these results affirm the effectiveness of these instruments in capturing the nuances of children’s mental health in these interactive settings during longitudinal CRI, showcasing the potential for integrating social robots in psychological assessments. Similar to our previous results in single-session CRI [81], the comparison between robot-administered and self-reported RCADS reveals significant distinctions that emphasize the potential benefits of integrating robotic assistance into mental health assessments. The analysis indicated that the robot-administered RCADS demonstrated a better fit in terms of construct validity, suggesting that interacting with a robot may facilitate a more coherent and reliable disclosure of symptoms from children. Moreover, in line with the literature, parents’ reports of RCADS also showed worse fit compared to the robot-administered version of RCADS, highlighting potential bias in parents’ reporting of their children’s mental health symptoms [85] compared to their children’s reporting of it to robots. As such, we can infer that robot-administered questionnaires can be considered reliable (See Section 4.1.6) and valid (See Section 4.2) and as such can be utilised as credible ground truth for the evaluation of mental wellbeing of children.
@@
Our results (Figure 3) also show that interacting with the robot online via Zoom or Skype has amplified the difference between the wellbeing clusters, especially the low tertile and high tertile. The effect is sustained over the course of three sessions as well as for all the sessions combined. We also observed a decrease in the RCADS ratings for the overall population over the course of three sessions. This could be a result of the regression to mean phenomenon which is common in longitudinal studies [86] and indicates that extreme observations have a tendency to move closer to the average value upon subsequent measurements[87]. The decrease in the RCADS ratings might also be attributed to the improvement in wellbeing stemming from the familiarity with a robot. Crossman et al [88] have demonstrated that interaction with a robot has increased the positive mood in children. Robots have also been used to alleviate anxiety during medical procedures [16]. Even in longitudinal HRI studies focussing on adult participants, interaction with robots has led to a decrease in stress levels [38, 53] and enhancement of participants’ mood [15].
@@
As seen from our findings (Figure 4), the robot-assisted mode of RCADS administration seems to be very effective in categorising the population across wellbeing clusters (especially low tertile and high tertile) as compared to traditional modes of self-reporting and parent-reporting of the children’s wellbeing. This could potentially be due to what prior studies have been reporting - i.e., that children consider robots as peers [89] and social companions [90] which could contribute to them being more candid about their emotions in front of a robot. Our finding is in line with our previous work [4] that also measured how wellbeing assessment of children differed across different modes of questionnaire administration, where the robot mode consisted of the children being physically co-located with the robot (i.e., a co-present robot).
@@
With regard to whether co-location affects wellbeing assessment, our findings suggest that there were no statistical differences found across the tertiles while comparing our remote study with our in-lab study presented in [4]. This is in line with previous research which suggests that co-location does not necessarily improve the objective of the CRI experience. Robinette et al [91] have demonstrated that there was minimal difference observed between the instruction-giving capabilities of a co-present robot, a remote robot and a virtual avatar robot. This is supported by Kennedy et al [46] who have shown that learning was not improved among children while interacting with the robot in person. Our study findings indicate that forthcoming robot-assisted initiatives for the assessment of wellbeing can be effectively conducted online, thus providing a logistical advantage to children to open up about their feelings from a comfortable environment (RQ1).
@@
With regard to the longitudinal changes in children’s perception of the robot (RQ2), we have observed (Figure 5) that most participants have maintained or improved their perception of the robot. This is in line with the previous literature [92] where prolonged exposure to robots has improved the opinions of Kindergarten children about the robot. Moreover, as seen from Figure 6, some children (low and med tertile) were more anxious in front of a physical robot while some children (high tertile) were more anxious in front of the remote robot. Prior research has indicated that people’s attitudes towards robots were governed by their prior knowledge about robots or even technology [93], with negative representation causing negative outlooks about robots [94]. We can also observe that children’s overall longitudinal perception (i.e. combined for all sessions) of the robot as a confidante as well as their anxiety towards the robot for all sessions combined does differ across varying levels of wellbeing (Figure 5). However, from our previous work [62], we have observed that the wellbeing of the children is independent of their anxiety towards the robot. This difference may arise from children’s varying perceptions of the robot that is physically present as compared to one that is remote, as observed from our findings (Figures 6). Additionally, prolonged exposure to robots might also alter children’s perception towards robots, potentially establishing a connection between their wellbeing and their attitude towards robots. Prior research [95, 96, 97] has indicated possible links between emotional and behavioural concerns and prolonged use of digital technology such as electronic gaming and the internet. Therefore, future online robot-assisted initiatives should consider participants’ outlooks, sentiments and existing/past emotional and behavioural concerns so that the technology can be utilised most effectively in the assessment of mental wellbeing of children.
@@
From our exploratory supplementary analyses regarding the effect of gender (RQ3) on the longitudinal robot-assisted evaluation of mental wellbeing of children, we have observed that the interaction with the online robot was able to effectively differentiate between the high tertiles and low tertiles across boys and girls for session 1 and only for boys for session 2. Strait et al [50] have observed that females respond favourably to changes in robot behaviour while males tend to prefer consistency. In our study, the experimental tasks used to evaluate wellbeing remained consistent across sessions, potentially leading to the statistically significant differences in wellbeing clusters for boys and not girls in session 2. Further, the effectiveness of the online robot mode of test administration (compared with co-present robot, self-report and parent-report responses) appears to be consistent across genders. Therefore, our study shows that the benefits of using online robot-assisted methods for assessing mental wellbeing hold true regardless of the participant’s gender. In addition, while the longitudinal perception of the online robot largely improved or remained consistent across genders, there was gender-related variability in how the robot was perceived as compared to our in-lab study presented in [62]. For instance, while comparing the responses of girls in the remote study and responses of girls in our in-lab study [62], we observed a decrease in the robot’s perception as a confidante (Figure 6(c)). We also observed a predominant increase in robot anxiety in boys (Figure 6(d)) between the remote study and our in-lab study presented in [62].
@@
In addition, when comparing responses to the perception of the role of the robot as a confidante of boys and girls in the remote study (Figure 5 (b)), we observed that girls had a higher opinion of the robot as compared to boys. The gender-based differences in user perception between the remote study and our in-lab study [62] could be attributed to the general difference in the attitudes across children with regard to technology [64]. Further, girls’ higher willingness to confide and share their emotions with others as compared to boys [13, 47]) could also contribute to their higher perception of the robot as a confidante in the remote study as compared to boys. Thus, in the design of future robot-led technologies for the assessment of mental wellbeing, it is crucial to take into account the differences in children’s perceptions of the robot and their willingness to open up, to enable an environment where they feel comfortable expressing their feelings truthfully.
@@
Moreover, in order to effectively aid the psychological support teams, it is important to consider how frequently should robotised assessments be conducted to gain insights into the children’s mental wellbeing. As we observed from our results in the remote study mentioned here and our in-lab study [4], one-off sessions, as well as longitudinal interactions with the robots, can both be effectively useful in understanding the mental wellbeing of children. From an HRI perspective, users’ responses, perceptions and engagement with regard to longitudinal interactions can be varied and must be taken into account while designing interactions. For example, studies have shown that prolonged exposure to robots helps participants promote more self-disclosure [15] and enables them to develop more positive opinions of the robot[92]. Studies have also shown that interaction with robots has resulted in improvement of mood as well as wellbeing of both children [88, 98] and adults [99, 15]. Thus, robot-driven assessments might contribute to enhancing the wellbeing of the participants, potentially affecting follow-up assessments in a positive manner. In contrast, studies have also shown that repeated interactions with the robot may also cause diminished engagement [39] and reduced interest to participate from the users [100]. Previous research has addressed this issue by implementing variations in the interaction experience such as changes in session activities [101] or adapting robot behaviour to align with user preferences to promote more social engagement [71].

### v3

@@ 4. Results
In summary, our results show that across all three sessions, RCADS responses have been significantly higher in the high tertile than in the low tertile. RCADS responses have also been significantly higher for the high tertile when comparing between modes of questionnaire administration (robot-report, self-report and parent-report).
Figure 5. (a) Longitudinal changes in children’s anxiety measurement attributed to the robot across all population groups (overall, girls and boys) categorised into wellbeing tertiles (low, med and high) using the SMFQ responses. (b) Gender-based variability in the role of the robot as a confidante in the remote study. (c) Total percentage of the population that had consistent or improved user perception. User perception changes across tertiles for all sessions combined for (d) perception of the robot as a confidante and (e) robot anxiety measurement. $*p<0.05$ corrected.
4.3. The effect of the longitudinal interactions and co-location on user’s perception of the robot (RQ2).
As seen in Figure 5, Friedman’s test indicated a statistically significant difference in the robot anxiety measurement scores (dependent variable) for the overall population group belonging to the low tertile, across the three sessions (independent variable) $\chi^{2}(2)=6.87,p=0.032$. Post hoc analysis with Wilcoxon signed-rank tests was conducted with a Bonferroni correction applied, resulting in statistical significance between the robot anxiety measurement responses for session 1 and session 3 ($Mdn_{session1}=16,Mdn_{session2}=14,p=0.041$). We have observed (Figure 5 (c)) that out of 40 participants (20 girls and 20 boys), most participants have either maintained or improved their perception of the robot. We defined improvement as an average increase in the self-reported ratings of the robot’s likeability, perceived intelligence, and role as a confidante, and an average decrease in the response ratings for the robot anxiety measurement i.e. the anxiety that the children attributed to the robot.
We did not find any statistically significant difference in the responses to the perception questionnaires (robot as confidante, robot anxiety measurement and likeability and perceived intelligence of the robot) across the tertiles throughout any of the sessions (session 1, session 2, and session 3). However, we have observed (Figure 5 (d)) that the Kruskal Wallis H test has shown a statistically significant difference ($H(2)=11.13,p=0.004$; $p=0.012$, Holm-Bonferroni corrected) in the children’s perception of the robot as a confidante (dependent variable) across tertiles (independent variable) for all the sessions combined. Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=33,Mdn_{high}=37,p=0.004$).
Kruskal Wallis H test has also shown a statistically significant difference ($H(2)=8.82,p=0.012$; $p=0.024$, Holm-Bonferroni corrected) in the robot anxiety measurement (dependent variable) across tertiles (independent variable) for all the sessions combined (Figure 5 (e)). Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=15,Mdn_{high}=21,p=0.012$). We did not find any statistical significance between tertiles for all sessions combined in the case of children’s likeability of the robot and their perceived intelligence of the robot.
As seen from Figure 6 (a), Wilcoxon rank sum test with a Bonferroni correction applied has indicated a statistically significant difference for the robot anxiety measurement in session 1 of the remote study and our in-lab study (Abbasi et al., 2023) for the med tertile of the overall population group ($Mdn_{remote}=18,Mdn_{in-lab}=24,p=0.038$). Further, Wilcoxon rank sum test (Figure 6 (b)) with a Bonferroni correction applied has also indicated a statistically significant difference for the perceived intelligence of the robot in session 1 of the remote study and our in-lab study (Abbasi et al., 2023) for the medium tertile of the overall population group ($Mdn_{remote}=24,Mdn_{in-lab}=21,p=0.045$). Wilcoxon rank sum test with a Bonferroni correction applied has also indicated a statistically significant difference between the perceived intelligence of the robot in session 1 of the remote study and our in-lab study (Abbasi et al., 2023) for the high tertile of the overall population group ($Mdn_{remote}=24,Mdn_{in-lab}=20.5,p=0.045$).
To sum up, our key findings indicate that longitudinally, robot anxiety measurement has been seen to decrease for the overall population. Further, the physical presence (remote vs. in-lab) of the robot does affect users’ anxiety attributed to the robot (robot anxiety increases for the low and med tertile but decreases for the high tertile) and the perceived intelligence of the robot (decreases across all tertiles).
Figure 6. User perception changes across the remote study and our in-lab study (Abbasi et al., 2023) categorised into wellbeing tertiles (low, med and high) using the SMFQ responses. $*p<0.05$ corrected.
4.4. Gender-based differences in the effectiveness of longitudinal wellbeing assessment and user perception (RQ3).
From our exploratory supplementary analyses, we have observed that in the case of girls (Figure 3 (a)), Kruskal Wallis H test has shown a statistically significant difference ($H(2)=10.64,p=0.005$; $p=0.020$, Holm-Bonferroni corrected) in the RCADS responses (dependent variable) for session 1 between the tertiles (independent variable). Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=6,Mdn_{high}=28,p=0.006$) and between med tertile and high tertile ($Mdn_{med}=6,Mdn_{high}=28,p=0.042$).
Further, Friedman’s test indicated a statistically significant difference (Figure 3(a)) in the RCADS scores (dependent variable) for girls belonging to the high tertile, across the three sessions (independent variable) ($\chi^{2}(2)=11.14,p=0.004$; $p=0.020$, Holm-Bonferroni corrected). Post hoc analysis with Wilcoxon signed-rank tests was conducted with a Bonferroni correction applied, resulting in statistical significance between the RCADS responses for session 1 and session 2 ($Mdn_{session1}=28,Mdn_{session2}=11.5,p=0.007$).
In the case of boys, as seen in Figure 3 (a), Kruskal Wallis H test has shown a statistically significant difference ($H(2)=7.91$, $p=0.019$, Holm-Bonferroni corrected) in the RCADS responses (dependent variable) for session 1 between the tertiles (independent variable). Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=9,Mdn_{high}=23,p=0.016$).
Kruskal Wallis H test has also shown a statistically significant difference ($H(2)=9.22,p=0.010$; $p=0.024$, Holm-Bonferroni corrected) in the RCADS responses (dependent variable) for session 2 between the tertiles (independent variable). Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=8.5,Mdn_{high}=20.5,p=0.009$).
We did not find any statistically significant difference longitudinally (between session 1, session 2, and session 3) between the RCADS responses for boys across all tertiles. We have also not observed any statistically significant differences (Figure 4) between the modes of RCADS administration (robot report, self-report, and parent-report) across all the tertiles for girls. In the case of boys (Figure 4) Wilcoxon rank sum test with a Bonferroni correction applied has indicated a statistically significant difference between the parent-report responses and the responses during session 1 with the robot for the high tertile ($Mdn_{parent}=7,Mdn_{robot}=23,p=0.013$). Further, even in the case of girls and boys, we have not found any statistically significant differences across the tertiles for the RCADS responses between session 1 for the remote study with the robot and the RCADS responses for our in-lab study presented in (Abbasi et al., 2022, 2024b).
We also did not find any statistical significance while comparing user perception responses to the questionnaires for girls or boys, across the tertiles, throughout any of the sessions. However, Wilcoxon rank sum test (Figure 5) has indicated a statistically significant difference between the children’s perception of the robot as a confidante for the high tertile between girls and boys ($Mdn_{girls}=37,Mdn_{boys}=33.5,W=67,p=0.012$; $p=0.030$, Holm-Bonferroni corrected). While considering user perception changes for girls (Figure 6(c)), Wilcoxon rank sum test with a Bonferroni correction applied has indicated a statistically significant difference for the perception of the robot as a confidante in session 1 of the remote study and our in-lab study (Abbasi et al., 2023) for the medium tertile of the overall population group ($Mdn_{remote}=37,Mdn_{in-lab}=31,p=0.038$). Further, in the case of boys (Figure 6(d)), Wilcoxon rank sum test with a Bonferroni correction applied has indicated a statistically significant difference for robot anxiety measurement in session 1 of the remote study and our in-lab study (Abbasi et al., 2023) for the med tertile of the overall population group ($Mdn_{remote}=20,Mdn_{in-lab}=24,p=0.048$).
In summary, across genders, we have observed that the RCADS responses were significantly higher for the high tertile for both boys and girls in session 1 and only for boys in session 2. We also observed that RCADS responses to the robot were significantly more negative for boys as compared to their parent’s perception of their wellbeing. Finally, girls have a higher perception of the robot as their confidante as compared with boys and the physical presence of the robot does affect this perception.
5. Discussion
The study explored the use of online robotised platform for the assessment of mental wellbeing of children where 40 children (20 girls and 20 boys, 8-13 years old) interacted with the Nao robot over Zoom or Skype for three sessions. Our findings indicate that online interactions with the robot can be used to effectively assess children’s wellbeing. Further, children’s perception of the robot either improved or stayed consistent and gender differences were also observed with girls viewing the robot more positively as a confidante.
@@ 5. Discussion
As seen from the construct validity analyses mentioned in Section 3.3, we employed a confirmatory PCA to explore the underlying structure of the SMFQ and the RCADS. We found that both scales predominantly revolve around a single principal component, suggesting a consistent underlying factor in each questionnaire explaining high variance of childrens’ well being. It is important to note that not all items in these questionnaires contributed equally to this primary factor. Specifically, some items in both the SMFQ and RCADS did not significantly load onto the principal component, indicating that they might not align as closely with the core factor that these scales measure within CRI. It should be acknowledged that RCADS (as used in this study) is aimed at assessing the concepts of generalised anxiety, panic disorder, and low mood (see Chorpita and et al., 2005), thereby potentially affecting factor loadings.
@@
Nevertheless, these results affirm the statistical effectiveness of these instruments in capturing the nuances of children’s mental health in these interactive settings during longitudinal CRI, showcasing the potential for integrating social robots in psychological assessments. Similar to our previous results in single-session CRI (Abbasi et al., 2024a), the comparison between robot-administered and self-reported RCADS reveals significant distinctions that emphasize the potential benefits of integrating robotic assistance into mental health assessments. This outcome represents a statistical validation rather than a behavioural indication. The analysis indicated that the robot-administered RCADS demonstrated a better fit in terms of construct validity, suggesting that interacting with a robot may facilitate disclosure of symptoms from children to assess their mental health in a statistically valid and reliable way. Moreover, in line with the literature, parents’ reports of RCADS also showed worse fit compared to the robot-administered version of RCADS, highlighting potential bias in parents’ reporting of their children’s mental health symptoms (Villatoro et al., 2018) compared to their children’s reporting of it to robots. As such, we can infer that robot-administered questionnaires can be considered reliable (See Section 3.2.6) and valid (See Section 3.3) and as such can be utilised as credible ground truth for the evaluation of mental wellbeing of children.
@@
Our results (Figure 3) also show that interacting with the robot online via Zoom or Skype has amplified the difference between the wellbeing clusters, especially the low tertile and high tertile. The effect is sustained over the course of three sessions as well as for all the sessions combined. We also observed a decrease in the RCADS ratings for the overall population over the course of three sessions. This could be a result of the regression to mean phenomenon which is common in longitudinal studies (Marsh and Hau, 2002) and indicates that extreme observations have a tendency to move closer to the average value upon subsequent measurements (Barnett et al., 2005). The decrease in the RCADS ratings might also be attributed to the improvement in wellbeing stemming from the familiarity with a robot. Crossman et al (Crossman et al., 2018) have demonstrated that interaction with a robot has increased the positive mood in children. Robots have also been used to alleviate anxiety during medical procedures (Rossi et al., 2022). Even in longitudinal HRI studies focussing on adult participants, interaction with robots has led to a decrease in stress levels (Russell et al., 2021; Laban et al., 2025) and enhancement of participants’ mood (Laban et al., 2024).
@@
As seen from our findings (Figure 4), the robot-assisted mode of RCADS administration seems to be very effective in categorising the population across wellbeing clusters (especially low tertile and high tertile) as compared to traditional modes of self-reporting and parent-reporting of the children’s wellbeing. This could potentially be due to what prior studies have been reporting - i.e., that children consider robots as peers (Kumar Singh et al., 2020) and social companions (Nalin et al., 2012b) which could contribute to them being more candid about their emotions in front of a robot. Our finding is in line with our previous work (Abbasi et al., 2022, 2024b) that also measured how wellbeing assessment of children differed across different modes of questionnaire administration, where the robot mode consisted of the children being physically co-located with the robot (i.e., a co-present robot).
@@
With regard to whether co-location affects wellbeing assessment, our findings suggest that there were no statistical differences found across the tertiles while comparing our remote study with our in-lab study presented in (Abbasi et al., 2022, 2024b). This is in line with previous research which suggests that co-location does not necessarily improve the objective of the CRI experience. Robinette et al (Robinette et al., 2016) have demonstrated that there was minimal difference observed between the instruction-giving capabilities of a co-present robot, a remote robot and a virtual avatar robot. This is supported by Kennedy et al (Kennedy et al., 2015) who have shown that learning was not improved among children while interacting with the robot in person. Our study findings indicate that forthcoming robot-assisted initiatives for the assessment of wellbeing can be effectively conducted online, thus providing a logistical advantage to children to open up about their feelings from a comfortable environment (RQ1).
@@
With regard to the longitudinal changes in children’s perception of the robot (RQ2), we have observed (Figure 5) that most participants have maintained or improved their perception of the robot. This is in line with the previous literature (Conti et al., 2019) where prolonged exposure to robots has improved the opinions of Kindergarten children about the robot. Moreover, as seen from Figure 6, some children (low and med tertile) were more anxious in front of a physical robot while some children (high tertile) were more anxious in front of the remote robot. This individual variability in attitudes could be due to their past experiences with robots and/or technology in general. Prior research has indicated that people’s attitudes towards robots were governed by their prior knowledge about robots or even technology (Sandoval et al., 2014), with negative representation causing negative outlooks about robots (Horstmann and Krämer, 2019). We can also observe that children’s overall longitudinal perception (i.e. combined for all sessions) of the robot as a confidante as well as their anxiety towards the robot for all sessions combined does differ across varying levels of wellbeing (Figure 5). However, from our previous work (Abbasi et al., 2023), we have observed that the wellbeing of the children is independent of their anxiety towards the robot. This difference may arise from children’s varying perceptions of the robot that is physically present as compared to one that is remote, as observed from our findings (Figures 6). Additionally, prolonged exposure to robots might also alter children’s perception towards robots, potentially establishing a connection between their wellbeing and their attitude towards robots. Prior research (Rikkers et al., 2016; Rapinda et al., 2021; Stenseng et al., 2020) has indicated possible links between emotional and behavioural concerns and prolonged use of digital technology such as electronic gaming and the internet. Therefore, future online robot-assisted initiatives should consider participants’ outlooks, sentiments and existing/past emotional and behavioural concerns so that the technology can be utilised most effectively in the assessment of mental wellbeing of children.
@@
From our exploratory supplementary analyses regarding the effect of gender (RQ3) on the longitudinal robot-assisted evaluation of mental wellbeing of children, we have observed that the interaction with the online robot was able to effectively differentiate between the high tertiles and low tertiles across boys and girls for session 1 and only for boys for session 2. Strait et al (Strait et al., 2015) have observed that females respond favourably to changes in robot behaviour while males tend to prefer consistency. In our study, the experimental tasks used to evaluate wellbeing remained consistent across sessions, potentially leading to the statistically significant differences in wellbeing clusters for boys and not girls in session 2. Further, the effectiveness of the online robot mode of test administration (compared with co-present robot, self-report and parent-report responses) appears to be consistent across genders. Therefore, our study shows that the benefits of using online robot-assisted methods for assessing mental wellbeing hold true regardless of the participant’s gender. In addition, while the longitudinal perception of the online robot largely improved or remained consistent across genders, there was gender-related variability in how the robot was perceived as compared to our in-lab study presented in (Abbasi et al., 2023). For instance, while comparing the responses of girls in the remote study and responses of girls in our in-lab study (Abbasi et al., 2023), we observed a decrease in the robot’s perception as a confidante (Figure 6(c)). We also observed a predominant increase in robot anxiety in boys (Figure 6(d)) between the remote study and our in-lab study presented in (Abbasi et al., 2023).
@@
In addition, when comparing responses to the perception of the role of the robot as a confidante of boys and girls in the remote study (Figure 5 (b)), we observed that girls had a higher opinion of the robot as compared to boys. The gender-based differences in user perception between the remote study and our in-lab study (Abbasi et al., 2023) could be attributed to the general difference in the attitudes across children with regard to technology (Jackson et al., 2008). Further, girls’ higher willingness to confide and share their emotions with others as compared to boys (Chandra and Minkovitz, 2006; Pearson, 2023)) could also contribute to their higher perception of the robot as a confidante in the remote study as compared to boys. Thus, in the design of future robot-led technologies for the assessment of mental wellbeing, it is crucial to take into account the differences in children’s perceptions of the robot and their willingness to open up, to enable an environment where they feel comfortable expressing their feelings truthfully.
@@
Moreover, in order to effectively aid the psychological support teams, it is important to consider how frequently should robotised assessments be conducted to gain insights into the children’s mental wellbeing. As we observed from our results in the remote study mentioned here and our in-lab study (Abbasi et al., 2022, 2024b), one-off sessions, as well as longitudinal interactions with the robots, can both be effectively useful in understanding the mental wellbeing of children. From an HRI perspective, users’ responses, perceptions and engagement with regard to longitudinal interactions can be varied and must be taken into account while designing interactions. For example, studies have shown that prolonged exposure to robots helps participants promote more self-disclosure (Laban et al., 2024) and enables them to develop more positive opinions of the robot(Conti et al., 2019). Studies have also shown that interaction with robots has resulted in improvement of mood as well as wellbeing of both children (Crossman et al., 2018; Leite et al., 2012) and adults (Bodala et al., 2021; Laban et al., 2024). Thus, robot-driven assessments might contribute to enhancing the wellbeing of the participants, potentially affecting follow-up assessments in a positive manner. In contrast, studies have also shown that repeated interactions with the robot may also cause diminished engagement (Serholt and Barendregt, 2016) and reduced interest to participate from the users (Leite et al., 2013). Previous research has addressed this issue by implementing variations in the interaction experience such as changes in session activities (Coninx et al., 2016) or adapting robot behaviour to align with user preferences to promote more social engagement (Ahmad et al., 2017).
In the context of our study, there were several key factors that influenced our methodology for using online-mediated interactions with social robots over other remote methods (such as chatbots, in-home robots or virtual animated characters of robots) that also aim to provide increased accessibility to the robots. Firstly, while chatbots have become increasingly popular in the mental health domain, often being used in mental health applications, they require some initiative from the participants (Huberty et al., 2021). This would be difficult to encourage in our study as the participants were children and could even lead to increased drop-out rates in the longitudinal study conducted here.
Secondly, in the case of in-home robots, there is a logistical disadvantage in terms of accessibility to every child  [^8] https://automationswitch.com/disadvantages-of-domestic-robots/, present in different parts of the world in comparison with smart devices required for online-mediated robots that are more readily available in households  [^9] https://explodingtopics.com/blog/smartphone-stats. As such, there would be monetary costs associated with purchasing either on the researchers’ end or on the participants’, creating inconvenience for each party.
Thirdly, in the case of virtual animated robot characters, prior research(Li, 2015) has suggested that they do not offer any notable advantage over online-mediated interaction with robots. Li et al (Li, 2015) have indicated that physical presence and not the physical embodiment of the robots have been shown to positively affect user responses. They demonstrated that co-present robots (present in the same room as the participants) have been shown to favourably influence users’ behaviour and preference in comparison with virtual (animated character of robots) or telepresent/online (physical robots present on a camera feed). In addition, Laban et al (Laban et al., 2024, 2025, 2022b) have shown that online-mediated interactions with robots (Pepper robot in their case) using videoconferencing platforms like Zoom have been shown to improve self-disclosure and mood of the participants. This finding aligns well with the familiarity and comfort that children developed with online platforms like Zoom and Skype over the Covid-19 pandemic, thus facilitating a more convenient and accessible interaction environment [^10] https://www.zoom.com/en/blog/how-zoom-supported-young-people-in-uk-during-pandemic/. Moreover, we also wanted to make valid comparisons with our previous works conducted in (Abbasi et al., 2022, 2024b). As such, we followed the experimental protocol conducted in (Abbasi et al., 2022, 2023, 2024b) and only modified the mode of interaction from in-person to online. Using a virtual robot (or agent) would have changed the robot’s form and appearance to an animated character making it difficult to ascertain if any changes in findings were due to the mode of the interaction (in-person vs. online) or the robot’s form (virtual animated character or the actual physical robot). Spitale et al (Spitale et al., 2023) have shown that in HRI for wellbeing studies changing robot form leads to significant changes in user perception and interaction. To this end, we employed robots that interacted with the participants through Zoom or Skype, to explore how they affect the assessment of mental wellbeing in children. In short, the aforementioned factors such as participants’ engagement, logistical accessibility, familiarity with online platforms and most importantly, methodological consistency to make valid comparisons with previous works (Abbasi et al., 2022, 2023, 2024b) influenced our approach to use online-mdeiated interactions with robots for the assessment of mental wellbeing of children.